OpenStudy (anonymous):

$8,000 is invested in two different accounts earning 3% and 4% interest. If the two accounts earn a total of $295 in interest, how much money was invested at 3%? $2,500 $3,500 $5,500

OpenStudy (anonymous):

try .03x + .04y = 295 x + y = 8000

OpenStudy (anonymous):

where x is the amount of money invested at 3% and y is the amount of money invested at 4%

OpenStudy (anonymous):

\[.03x+.04(8000-x)=295\] is a start

OpenStudy (anonymous):

mutiply by 100 to get rid of annoying decimals get \[3x+4(8000-x)=29500\] and solve that one

OpenStudy (anonymous):

so its 2,500

OpenStudy (anonymous):

yes